Internet/Intranet Connection Downtime

Who/Where: 
College of Arts and Sciences
When: 
Friday, April 19, 2013 - 08:00 to Tuesday, April 23, 2013 - 09:45
Why: 
Fiber Optic Problem
Account: 

* Upon FO testing yesterday (From CC to NEDA Padmount, NEDA Padmount to CAS Oblation, CAS Oblation to RH Hallway, RH Hallway to CAS Library) using Visual Fault Locator with the help of Junjun, we were able to detect defective FO patch cord connected to the CAS-RH Switch located at the CAS Library.
* The FO patch cable has been replaced yesterday afternoon but still the connection between our core switch and CAS-RH switch is not present, so we decided to track down the FO cores from the RH Oblation to the NEDA Padmount again. Tests were successful.
* For the 2nd time, we test the cores from Padmount at NEDA to Computer Center, we found one defective core and we successfully found another as replacement.
* Around 5PM yesterday with the help of Sir Ricky from CAS Library, I conducted another test from Computer Center to CAS Library and the output signal of the VFL was very poor on CAS side. (FO Problem still exists)
* Early this morning, Junjun and I conducted a test again at NEDA Padmount using VFL and media converter:
--> First, we removed the cores going to CAS and put it on the media converter, the link was okay, defective CAS FO link has been ruled out.
--> Second, we removed the connection from Computer Center and put it on media converter, the link was okay.
--> Then, we put back the cores from CC to the coupler and used a FO patch cable directly to the media converter, but there is no connection. Even when we used a different coupler
--> Using VFL we traced the FO patch cable and the result was okay.
--> And then we decided to use 2 different cores from CC to NEDA Padmount. The connection of CAS has been re-established.